(D2 - 4) y= xe2 , xex is yp = (ax + b) ex

1
Expert's answer
2021-06-08T04:37:31-0400

"k^2-4=0"

"k_{1}=-2,k_2=2"


"y=c_1e^{-2x}+c_2e^{2x}"


"y_p=(ax^2+bx)e^{2x}"

"y'_p=(2ax+b)e^{2x}+2(ax^2+bx)e^{2x}=(2ax^2+(2a+2b)x+b)e^{2x}"

"y''_p=(4ax+2a+2b)e^{2x}+2(2ax^2+(2a+2b)x+b)e^{2x}"

"4ax^2+8ax+4bx+2a+4b-4ax^2-4bx=x"

"a=1\/8,b=-1\/16"


"y=c_1e^{-2x}+c_2e^{2x}+(x^2\/8-x\/16)e^{2x}"